# Beyond Static AI Evaluations: Advancing Human Interaction Evaluations for LLM Harms and Risks

### v2

@@ 5 Example Evaluations
1 Identifying risk and/or harm Research question LLMs have been shown to produce political messages that rival the persuasiveness of human messages, even those written by expert political consultants (Hackenburg et al. 2023). The majority of research on such persuasion risks has focused on attitude change due to exposure to persuasive LLM output. It is equally important to study the mechanisms of uplift in message persuasiveness when users access an LLM to co-write a persuasive message. This can aid in understanding LLMs’ persuasive capabilities and their limitations (Phuong et al. 2024). Thus, here, we ask: how does the use of an LLM lead to a change in human writing performance — quality and persuasiveness — of political messaging in opinion pieces?
@@ 6.1 Limitations
Our framework and taxonomy are starting points for safety-focused HIEs and are focused on evaluation units consisting of one human and one model. More detailed taxonomizing is necessary as models continue to be integrated in real-world applications and new interaction forms involving multiple human actors and models emerge (Li et al. 2023; Ward et al. 2024). Additionally, our framework does not fully consider the effects of feedback loops in human-LLM interaction which complicate discerning whether outcomes are driven by the model’s influence on humans or human input to the model. Further research into the conceptual and empirical risks of these loops is needed (Pan et al. 2024; Krauth, Wang, and Jordan 2022).
@@ 6.2 Recommendations
  • Costs: experiments, particularly long experiments or powered experiments with large sample sizes, can require significant funds to compensate participants and develop and use online platforms.
@@
This paper serves as a step towards organizing and sharing design-level considerations for safety-focused HIEs. We recommend further developing accessible protocols, guides, and standardized test suites, similar to those available for non-interactive evaluations, to specifically facilitate human participation in model evaluation (UK AI Safety Institute 2024; METR 2023). An example of this is CheckMate, a prototype platform which records interaction traces of humans interacting with LLMs (Collins et al. 2023). Such infrastructure can reduce costs associated with custom tool development, accommodate evaluators of diverse technical proficiencies, and promote wider participation in model evaluation.
@@ Appendix A Appendix
Hackenburg et al. (2023) persuasion experiments Y marginal exposure 1

### v4

@@ 5 Example Evaluations
1 Identifying risk and/or harm Research question LLMs have been shown to produce political messages that rival the persuasiveness of human messages, even those written by expert political consultants (hackenburgcomparing). The majority of research on such persuasion risks has focused on attitude change due to exposure to persuasive LLM output. It is equally important to study the mechanisms of uplift in message persuasiveness when users access an LLM to co-write a persuasive message. This can aid in understanding LLMs’ persuasive capabilities and their limitations (Phuong et al. 2024). Thus, here, we ask: how does the use of an LLM lead to a change in human writing performance — quality and persuasiveness — of political messaging in opinion pieces?
@@ 6.1 Limitations
Our framework and taxonomy are starting points for safety-focused HIEs and are focused on controlled, human subject experiments and evaluation units consisting of one human and one model. More detailed taxonomizing is necessary as models continue to be integrated in real-world applications and new interaction forms involving multiple human actors and models emerge (Li et al. 2023; Ward et al. 2024). Additionally, our framework does not fully consider the characteristics and effects of feedback loops in human-LLM interaction which complicate discerning whether outcomes are driven by the model’s influence on humans or human input to the model. Further research into the conceptual and empirical risks of these loops is needed (Pan et al. 2024; Krauth, Wang, and Jordan 2022).
@@ 6.2 Recommendations
  • Costs and scale: experiments, particularly long experiments or powered experiments with large sample sizes, can require significant funds to compensate participants and develop and use online platforms, and thus are difficult to scale.
@@
This paper serves as a step towards organizing and sharing design-level considerations for safety-focused HIEs. We recommend further developing accessible protocols, guides, and standardized test suites, similar to those available for non-interactive evaluations, to specifically facilitate human participation in model evaluation (UK AI Safety Institute 2024; METR 2023). An example of this is CheckMate, a prototype platform which records interaction traces of humans interacting with LLMs (Collins et al. 2023). Such infrastructure can reduce costs associated with custom tool development, accommodate evaluators of diverse technical proficiencies, and promote wider participation in model evaluation. Further research should also investigate methods to create extensive automated evaluations informed by scoped HIEs.
@@ Appendix A Appendix
hackenburgcomparing persuasion experiments Y marginal exposure 1